OpenStudy (anonymous):

What is csc(5pi/3)+tan(5pi/6)+sec(7pi/4)?

OpenStudy (e.mccormick):

\[\csc \frac{5\pi}{3} + \tan \frac{5\pi}{6}+ \sec \frac{7\pi}{4}\] Each of those has a value on the unit circle. Find those values and add them. That is all they are asking you to do.

OpenStudy (anonymous):

This is just a test; my first question was a glitch; thank you anyway, i gave you a medal :)

OpenStudy (e.mccormick):

Ah, hehe. Well, don't have too much fun. The system generally works and if you learn a little LaTeX or use the equation editor you can make some pretty nice looking calculations.

OpenStudy (anonymous):

do you have a aops account

OpenStudy (anonymous):

just asking

OpenStudy (e.mccormick):

Nope.